# LightGaussian: Unbounded 3D Gaussian Compression with 15x Reduction and 200+ FPS

### 07 Dec 2023 v3

@@ 2 Related Works
Neural radiance field (NeRF) [44] uses a multi-layer perceptron (MLP) to represent a scene, and this compact representation has brought view synthesis quality to a new stage. However, NeRF suffers from extremely slow inference challenges. Follow-ups either explore ray re-parameterizations[4, 5], explicit spatial data structures [68, 21, 40, 31, 56, 46, 10], caching and distillation [61, 22, 28, 49], or ray-based representations [2, 55] for speeding up. Still, NeRF-based methods struggle to achieve real-time rendering speed in practical large-scale scenes, which is caused by the multiple queries needed for rendering a single pixel, limiting their practical use. Point-based representations (e.g., Gaussian) have been extensively explored in various domains, including shape reconstruction [33], molecular structures modeling [7], and the replacement of point-cloud data [15]. Additionally, these representations are utilized in shadow [47] and cloud rendering applications [42]. Recently, 3D Gaussian Splatting (3D-GS) [32], combining the idea of point-based rendering and splatting techniques for rendering, achieves real-time speed with comparable rendering quality to the best MLP-based renderer, Mip-NeRF 360 [5]. Although promising, the heavy storage requirement to store all attributes attached to the Gaussians often requires gigabyte-level disk space for saving a single unbounded scene, and millions of Gaussians hinder the rendering efficiency of 3D-GS.
@@ 3.1 Background: 3D Gaussian Splatting
Here, $c_{i}$, $\alpha_{i}$ represents the color and opacity of this point computed by a Gaussian with covariance $\boldsymbol{\Sigma}$ multiplied by an optimizable per-point opacity and SH color coefficients. In summary, each Gaussian point is characterized by attributes including: position $\boldsymbol{X}\in\mathbb{R}^{3}$, color defined by spherical harmonics coefficients $\boldsymbol{C}\in\mathbb{R}^{(k+1)^{2}}\times 3$ (where $k$ represents the degrees of freedom), opacity $\alpha\in\mathbb{R}$, rotation factor $\boldsymbol{r}\in\mathbb{R}^{4}$, and scaling factor $\boldsymbol{s}\in\mathbb{R}^{3}$.
@@ 3.2 Gaussian Pruning & Recovery
Gaussian densification [32], which involves cloning and splitting the initial SfM point cloud, is employed to address the challenge of insufficient coverage and is used to model small-scale geometry as well as detailed scene appearance. While this strategy leads to significantly improved reconstruction quality, it results in the number of Gaussians growing from thousands to millions after optimization. Such an explicit point-based representation with a large number of Gaussians requires an extremely significant storage overhead. However, pruning the Gaussians based on simplistic criteria (e.g., point opacity) can lead to a substantial degradation in modeling performance, especially where the intricate scene structure may be eliminated, as is demonstrated in Fig. 3.
Figure 3: Zero-shot Opacity-based Pruning. A significant number of Gaussians exhibit small opacity values (top). Simply utilizing Gaussian opacity as an indicator for pruning the least important Gaussians results in the rendered image losing intricate details (bottom), with the PSNR dropping from 27.2 to 25.3. This has inspired us to find better criteria to measure global significance in terms of rendering quality.
Drawing inspiration from the success of representative neural network pruning techniques [26], which eliminate less impactful neurons without compromising the network’s overall performance, we tailor a general pruning paradigm for point-based representation to reduce the over-parameterized point number in a manner that preserves the original accuracy. Therefore, identifying the most representative redundant Gaussians with recoverable accuracy is a crucial step in our approach.
@@
Simply relying on Gaussian opacity as a significance criterion leads to suboptimal Gaussian pruning, prompting the need for a more effective formula. Inspired by Equation 3, these 3D Gaussians can be rasterized by projecting them onto a specific camera viewpoint for image rendering. The significance of each Gaussian can then be quantified based on its contribution to each pixel across all training views, with similar principle in magnitude network pruning [38]. Consequently, we iterate over all training pixels to calculate the hit count of each Gaussian, factoring in both Gaussian opacity and volume:
$\text{GS}_{\text{j}}$ $=\sum_{i=1}^{MHW}\mathbbm{1}(\text{G}(\boldsymbol{X}_{j}),r_{i})\cdot\sigma_{j}\cdot\gamma_{j}(\boldsymbol{\Sigma}),$ (4)
@@
The use of absolute Gaussian volume tends to exaggerate the importance of background Gaussians, leading to the immoderate pruning of Gaussians that model intricate geometry. Therefore, we propose a more adaptive way to measure the dimension of its volume:
@@
We rank all the Gaussians based on the computed global significance scores, and provide a quantitative basis for pruning Gaussians with the lowest scores. It is worth noting that an overly aggressive pruning ratio can be detrimental to performance. Stripping away too many Gaussians can lead to a noticeable degradation in the model’s accuracy and visual fidelity since incomplete scene coverage. To mitigate this, a phase of short Gaussian co-adaptation is essential, which involves a joint adjustment of the Gaussians’ attributes, allowing them to adapt and compensate for the loss incurred by pruning.
@@ 3.3 Distilling into Compact SHs
To achieve a balance between disk storage efficiency and scene reflectance quality, we propose transferring knowledge from well-trained high-degree SHs to their compact counterparts (lower-degree) via data distillation:
$\mathcal{L}_{\text{distill}}(\text{G}_{\text{SH2-deg}})=\frac{1}{N}\sum_{i=1}^{N}\left\lVert\boldsymbol{C}_{\text{teacher}}-\boldsymbol{C}_{\text{student}}\right\rVert_{2}^{2}.$ (6)
@@ 3.4 VecTree Attribute Compression
Vector quantization is well-explored in voxel-based NeRF representations [39, 68, 57, 72]. Although a high compression rate can be reached, Gaussian attributes cannot be easily quantized in the same fashion. The reason is that each attribute of the explicit representation has its own physical meaning, especially when representing large-scale 3D scenes, and attributes such as opacity, rotation, and scale are more sensitive to such discretized representation, with a dramatic drop in accuracy. Therefore, we propose two ways to mitigate this gap: first, we reutilize the significance score from Sec. 3.2 and perform excessive vector quantization on the SHs of the least important Gaussians. For the Gaussian location, we draw from point cloud compression and adopt octree-based lossless compression into our framework. For the remaining important SHs and other attributes representing the Gaussian shape, rotation, and opacity, we save these features in float16 format.
@@
The codebook under vector quantization (VQ) is initialized with $K$ codes within it, denoted as $\boldsymbol{C}={\boldsymbol{c}_{1},\boldsymbol{c}_{2},\boldsymbol{c}_{3},\boldsymbol{c}_{K}}$ to represent the Gaussian sets $\boldsymbol{G}={\boldsymbol{g}_{1},\boldsymbol{g}_{2},\boldsymbol{g}_{3},\boldsymbol{g}_{N}}$ where $K\ll N$. To assign each code to the Gaussian feature, we randomly select a batch of Gaussians at each iteration and calculate the Euclidean distance between the selected Gaussian and each code vector in the codebook to determine which code the Gaussian is associated with. Then, a significance-weighted codebook optimization is applied: $\mathbf{c}_{i}:=\lambda\cdot\mathbf{c}_{i}+(1-\lambda)\cdot\sum{\mathbf{g}_{j}\in\mathcal{R}(\mathbf{c}_{k})}\text{GS}_{j}\cdot\mathbf{g}_{j}$, where $\mathbf{g}_{j}\in\mathcal{R}(\mathbf{c}_{i})$ denotes the $j$-th Gaussian assigned to the $i$-th code, and $\lambda$ controls the contribution from the moving average. We only perform excessive VQ on the least 60% of Gaussians, since we are representing complex, large-scale scenes.
@@
Unstructured and high-precision 3D points pose a challenge for efficient storage, especially for large-scale scenes. Applying lossy compression to Gaussian locations shows significant rendering degradation, as the Gaussian location is sensitive to the subsequent rasterization accuracy. Therefore, drawing from the Point Cloud Compression literature [53], which applies lossless encoding on the Gaussian location, utilizing the octree geometry codec in G-PCC [45] consumes only a very small amount of bits. Specifically, let $(X_{i}=(x_{i},y_{i},z_{i}))_{i=1\ldots N}$ be the set of 3D positions associated with the points of the input point cloud. The L-PCC encoder [53] computes the quantized positions $(\hat{X}_{i})_{i=1...N}$ as follows: $\hat{X}_{i}=\left\lfloor(X_{i}-X_{shift})\times s\right\rfloor,$ where $X_{shift}$ and $s$ are pre-defined parameters and we set all octree-related parameters the same as PCGCv2 [62].
For the decoding stage, the reconstructed positions $(\hat{X}_{i})_{i=1}^{N}$ are generated by applying the following inverse quantization process: $\hat{X}_{i}=\frac{\hat{X}i}{s}+X_{shift}$.
@@ 4.2 Experimental Results
We compare the rendering result of all adopted baselines, with the highlight on intricate details and the background regions in Fig. 4. We can observe Mip-NeRF360, 3D-GS and our LightGaussian achieves the best visual quality, whereas Mip-NeRF360 requires more than 30 seconds to render an image of resolution 1080p, 3D-GS requires 17$\times$ times more disk storage requirement.
@@ 6 Acknowledgements
Tables A6, A7, and A8 present the comprehensive error metrics compiled for our evaluation across all real-world scenes (Mip-NeRF360 and Tank and Temple datasets). Our method not only compresses the average model size from 727MB to 42MB, but also consistently demonstrates comparable metrics with 3D-GS on all scenes. LightGaussian additionally shows better rendering quality than Plenoxel, INGP, mip-NeRF360, and VQ-DVGO.

### 06 Feb 2024 v4

@@ 2 Related Works
Neural radiance field (NeRF) [43] uses a multi-layer perceptron (MLP) to represent a scene, and this compact representation has brought view synthesis quality to a new stage. However, NeRF suffers from extremely slow inference challenges. Follow-ups either explore ray re-parameterizations[4, 5], explicit spatial data structures [66, 21, 39, 31, 55, 45, 10], caching and distillation [60, 22, 28, 48], or ray-based representations [2, 54] for speeding up. Still, NeRF-based methods struggle to achieve real-time rendering speed in practical large-scale scenes, which is caused by the multiple queries needed for rendering a single pixel, limiting their practical use. Point-based representations (e.g., Gaussian) have been extensively explored in various domains, including shape reconstruction [33], molecular structures modeling [7], and the replacement of point-cloud data [15]. Additionally, these representations are utilized in shadow [46] and cloud rendering applications [41]. Recently, 3D Gaussian Splatting (3D-GS) [32], combining the idea of point-based rendering and splatting techniques [72] for rendering, achieves real-time speed with comparable rendering quality to the best MLP-based renderer, Mip-NeRF 360 [5]. Although promising, the heavy storage requirement to store all attributes attached to the Gaussians often requires gigabyte-level disk space for saving a single unbounded scene, and millions of Gaussians hinder the rendering efficiency of 3D-GS.
@@ 3.1 Background: 3D Gaussian Splatting
Here, $c_{i}$, $\alpha_{i}$ represents the color and opacity of this point computed by a Gaussian with covariance $\boldsymbol{\Sigma}$ multiplied by an optimizable per-point opacity and SH color coefficients. In summary, each Gaussian point is characterized by attributes including: position $\boldsymbol{X}\in\mathbb{R}^{3}$, color defined by spherical harmonics coefficients $\boldsymbol{C}\in\mathbb{R}^{(k+1)^{2}}\times 3$ (where $k$ represents the degrees of freedom), opacity $\alpha\in\mathbb{R}$, rotation factor $\mathbf{R}\in\mathbb{R}^{4}$, and scaling factor $\mathbf{S}\in\mathbb{R}^{3}$.
@@ 3.2 Gaussian Pruning & Recovery
Gaussian densification [32], which involves cloning and splitting the initial SfM point cloud, is employed to address the challenge of insufficient coverage and is used to model small-scale geometry as well as detailed scene appearance. While this strategy leads to significantly improved reconstruction quality, it results in the number of Gaussians growing from thousands to millions after optimization. Such an explicit point-based representation with a large number of Gaussians requires an extremely significant storage overhead. Drawing inspiration from the success of representative neural network pruning techniques [26], which eliminate less impactful neurons without compromising the network’s overall performance, we tailor a general pruning paradigm for point-based representation to reduce the over-parameterized point number in a manner that preserves the original accuracy. Therefore, identifying the most representative redundant Gaussians with recoverable accuracy is a crucial step in our approach. However, pruning the Gaussians based on simplistic criteria (e.g., point opacity) can lead to a substantial degradation in modeling performance, especially where the intricate scene structure may be eliminated, as is demonstrated in Fig. 3.
Figure 3: Zero-shot Opacity-based Pruning. A significant number of Gaussians exhibit small opacity values (top). Simply utilizing Gaussian opacity as an indicator for pruning the least important Gaussians results in the rendered image losing intricate details (bottom), with the PSNR dropping from 27.2 to 25.3. This has inspired us to find better criteria to measure global significance in terms of rendering quality. The accumulated PDF is equal to 1.
@@
Simply relying on Gaussian opacity as a significance criterion leads to suboptimal Gaussian pruning, prompting the need for a more effective formula. Inspired by Eq 3, these 3D Gaussians can be rasterized by projecting them onto a specific camera viewpoint for image rendering. The initial significance score of each Gaussian can then be quantified based on its contribution to each pixel (a.k.a rays, $r_{i}$) across all training views using the criteria $\mathbbm{1}(\text{G}(\boldsymbol{X}_{j}),r_{i})$. Consequently, we iterate over all training pixels to calculate the hit count of each Gaussian. The score is further refined by the adjusted 3D Gaussian’s volume $\large{\gamma}$$(\boldsymbol{\Sigma}_{j})$ and opacity $\sigma_{j}$. The volume calculation equation is V${(\boldsymbol{\Sigma}_{j})}=\frac{4}{3}\pi abc$, where $abc$ are the 3 dimensions of Scale ($\mathbf{S}$). We implement this in PyTorch as torch.prod(gaussians.scaling, dim=1) without considering the constant value $\frac{4}{3}\pi$. Overall, the global significant score can be summarized as:
$\text{GS}_{\text{j}}$ $=\sum_{i=1}^{MHW}\mathbbm{1}(\text{G}(\boldsymbol{X}_{j}),r_{i})\cdot\sigma_{j}\cdot{\large{\gamma}}(\boldsymbol{\Sigma}_{j}),$ (4)
@@
However, the use of Gaussian volume tends to exaggerate the importance of background Gaussians, leading to the immoderate pruning of Gaussians that model intricate geometry. Therefore, we propose a more adaptive way to measure the dimension of its volume:
@@
We rank all the Gaussians based on the computed global significance scores, and provide a quantitative basis for pruning Gaussians with the lowest scores. It is worth noting that an overly aggressive pruning ratio can be detrimental to performance. Stripping away too many Gaussians can lead to a noticeable degradation in the model’s accuracy and visual fidelity since incomplete scene coverage. To mitigate this, we propose to utilize photometric loss on original training views to optimize (adapt) the remaining Gaussians, which jointly adjusts their attributes over 5k steps, without further densification, which can compensate for the loss incurred by pruning.
@@ 3.3 Distilling into Compact SHs
To achieve a balance between model size and scene reflectance quality, we propose transferring knowledge from well-trained high-degree SHs to their compact counterparts (lower-degree) via distillation. Specifically, we distill the teacher model (with full-degree SHs, and is frozen) to the student model (with truncated lower-degree SHs). The supervisions are from the difference of the predicted pixel intensities $\boldsymbol{C}$ from the two models :
$\mathcal{L}_{\text{distill}}=\frac{1}{HW}\sum_{i=1}^{HW}\left\lVert\boldsymbol{C}_{\text{teacher}}(r_{i})-\boldsymbol{C}_{\text{student}}(r_{i})\right\rVert_{2}^{2}.$ (6)
@@ 3.4 VecTree Attribute Compression
Vector quantization is well-explored in voxel-based NeRF representations [38, 66, 56, 70]. Although a high compression rate can be reached, Gaussian attributes cannot be easily quantized in the same fashion. The reason is that each attribute of the explicit representation has its own physical meaning and complex correlation (e.g., the rotation and scaling formulate the covariance matrix), and we empirically find that certain attributes such as opacity, rotation, and scale are more sensitive to such discretized representation, with a dramatic drop in accuracy. Therefore, we propose two ways to mitigate this gap: first, we reutilize the significance score from Sec. 3.2 and perform excessive vector quantization on the SHs of the least important Gaussians. For the Gaussian location, we draw from point cloud compression and adopt octree-based lossless compression into our framework. For the remaining important SHs and other attributes representing the Gaussian shape, rotation, and opacity, we save these features in float16 format.
@@
VQ partitions the Gaussians $\mathcal{G}=\{\mathbf{g}_{1},\mathbf{g}_{2},\ldots,\mathbf{g}_{N}\}$ to the K codes in the codebook $\mathcal{C}=\{\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{K}\}$, where each $\mathbf{g}_{j},\mathbf{c}_{k}\in\mathbb{R}^{d}$ and K $\ll$ N. $d$ means attributes’ dimension. Specifically, $\mathcal{C}$ is initialized via K-means. Then, VQ iteratively samples a batch of $\mathcal{G}$, associates them to the closest codes by euclidean distance, and updates each $\mathbf{c}_{k}$ via moving average rule: $\mathbf{c}_{k}=0.8\cdot\mathbf{c}_{k}+(1-0.8)\cdot 1/\mathbf{T}_{k}\cdot\sum_{\mathbf{g}_{j}\in\mathcal{R}(\mathbf{c}_{k})}\mathbf{GS}_{j}\cdot\mathbf{g}_{j}$, where $\mathbf{T}_{k}=\sum_{\mathbf{g}_{j}\in\mathcal{R}(\mathbf{c}_{k})}{\mathbf{GS}_{j}}$ is the total significance score assigned to the code vector $\mathbf{c}_{k}$, $\mathcal{R}(\mathbf{c}_{k})$ is the set of Gaussians associated to the $k$-th code.
@@
To deal with the unstructured and high-precision 3D points location, we avoid the lossy compression which is sensitive to the subsequent rasterization accuracy. Drawing inspiration from the Point Cloud Compression literature [52], which applies lossless encoding on the Gaussian location, utilizing the octree geometry codec in G-PCC [44] consumes only a very small amount of bits.
@@ 4.2 Experimental Results
We compared the rendering results between 3D-GS and ours, with an emphasis on intricate details and background regions as illustrated in Fig. 4, to highlight the differences. It is observable that 3D-GS and our LightGaussian achieve similar visual quality, even in thin structures within the scenes (the first and third rows). The Gaussian compression method slightly affects the specular effect on the surface (see the second and the fourth rows).
@@ 4.3 Ablation Studies
Degradation & Speed vs. Compression Ratio?
To strike a balance among model compression ratio and the reduced model size as well as the rendering speed, we study the effect of various ratios for Gaussian Pruning & Recovery, SHs Distilling, and VecTree Compression. As can be seen in the Fig 7 and Tab 5. The ratios selected represent a balanced point to optimize the trade-off between model size, FPS, and rendering quality (see Tab 5).
Figure 7: Visualization of The effect from different prune and VQ ratio.
Exp ID Model setting Model Size $\downarrow$ FPS $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
[1] 3D-GS 353.27 MB 192 31.687 0.927 0.200
[2] 3D-GS (w/ 40% pruned Gaussian) 211.96 MB 245 31.930 0.928 0.198
[3] [2] + 2-degree SH compactness 140.16 MB 238 30.625 0.917 0.208
[4] [3] + proposed distillation 140.16 MB 243 31.754 0.926 0.201
[5] [2] + 1-degree SH compactness 88.89 MB 249 29.173 0.900 0.223
[6] [5] + proposed distillation 88.89 MB 244 31.440 0.923 0.205
[7] [4] + 55% vector quantization 42.92MB 244 31.626 0.924 0.204
Table 5: Ablation under a mild compression ratio: With a Gaussian pruning ratio of 40% and VQ ratio of 55%, LightGaussian nearly preserves the rendering quality (SSIM decrease from 0.927 to 0.924), while achieving an $\sim$87reduction in model size and enhancing the FPS from 192 to 244.
@@ 6 Acknowledgements
Tables A7, A8, and A9 present the comprehensive error metrics compiled for our evaluation across all real-world scenes (Mip-NeRF360 and Tank and Temple datasets). Our method not only compresses the average model size from 727MB to 42MB, but also consistently demonstrates comparable metrics with 3D-GS on all scenes. LightGaussian additionally shows better rendering quality than Plenoxel, INGP, mip-NeRF360, and VQ-DVGO.